# Correlated electronic structures and unconventional superconductivity in bilayer nickelate heterostructures

### v1

@@ Discussion
Figure S5: | Filling $n$-dependence of $A({\bf k},\omega)$ (top panels) and FS (lower panels) calculated at $U=$3.5 eV, $J=$0.56. In all panels, $U=3.5$ eV, $J$=0.56 eV, $T=200K$.
Figure S6: | Temperature dependence of $A({\bf k},\omega)$ (top panels) and FS (lower panels) calculated in CDMFT at $U=3.5$ eV, $J$=0.56 eV, $n$=1.3. a-b $T$=200 K; c-d $T$=50K.
Figure S7: The spin susceptibilities and corresponding gap functions on the FS. a, d $n=1.25$, $U=3.5$ eV and $U_{\textrm{eff}}=0.205$eV. b, e $n=1.3$, $U=3.77$ eV and $U_{\textrm{eff}}=0.14$eV. c, f $n=1.3$, $U=3$ eV and $U_{\textrm{eff}}=0.31$eV.
Figure S8: The dependence of $\lambda$ on $U_{\textrm{eff}}$ for a $U=3.77$ eV and b $U=3$ eV. The filling is $n=1.3$.

### v2

@@ Discussion
Figure S5: | Filling $n$-dependence of $A({\bf k},\omega)$ (top panels) and FS (lower panels) calculated at $U=$3.6 eV, $J=$0.56. In all panels, $U=3.6$ eV, $J$=0.56 eV, $T\sim$106K.
Figure S6: | Temperature dependence of $A({\bf k},\omega)$ (top panels) and FS (lower panels) calculated in CDMFT at $U=3.6$ eV, $J$=0.56 eV, $n$=1.3. a-b $T$=200 K; c-d $T$=50K.
Figure S7: | The spin susceptibilities and corresponding gap functions on the FS. a, b $n=1.25$, $U=3.6$ eV and $U_{\textrm{eff}}=0.17$eV. c, d $n=1.3$, $U=3.77$ eV and $U_{\textrm{eff}}=0.14$eV. e, f $n=1.3$, $U=3.6$ eV and $U_{\textrm{eff}}=0.205$eV. g, h $n=1.3$, $U=3$ eV and $U_{\textrm{eff}}=0.312$eV.
Figure S8: | The dependence of $\lambda$ on $U_{\textrm{eff}}$. a $U=3.77$ eV. b $U=3.6$ eV. c $U=3$ eV. The filling is $n=1.3$.